OpenStudy (anonymous):

if a certain number is added to the numerator and denom of 7/9, the result is 5/6. find the #'s

OpenStudy (phi):

call the number x. this problem says \[ \frac{7+x}{9+x}= \frac{5}{6} \] you solve this by multiplying both sides by (9+x)6 can you do that?

OpenStudy (anonymous):

no sure I am understanding